# Spurious-Aware Prototype Refinement for Reliable Out-of-Distribution Detection

### v1

@@ Abstract
Out-of-distribution (OOD) detection is crucial for ensuring the reliability and safety of machine learning models in real-world applications, where they frequently face data distributions unseen during training. Despite progress, existing methods are often vulnerable to spurious correlations that mislead models and compromise robustness. To address this, we propose SPROD, a novel prototype-based OOD detection approach that explicitly addresses the challenge posed by unknown spurious correlations. Our post-hoc method refines class prototypes to mitigate bias from spurious features without additional data or hyperparameter tuning, and is broadly applicable across diverse backbones and OOD detection settings. We conduct a comprehensive spurious correlation OOD detection benchmarking, comparing our method against existing approaches and demonstrating its superior performance across challenging OOD datasets, such as CelebA, Waterbirds, UrbanCars, Spurious Imagenet, and the newly introduced Animals MetaCoCo. On average, SPROD improves AUROC by 4.7% and FPR@95 by 9.3% over the second best.
@@ 1 Introduction
Machine learning systems in real-world applications often encounter out-of-distribution (OOD) inputs, which are samples from distributions different from the training data. These inputs require cautious handling to prevent incorrect overconfident predictions during inference [1]. This makes OOD detection crucial, as it aims to identify whether an input belongs to the known distribution or not. Yet, deep neural networks, widely used in vision tasks [2, 3, 4, 5], tend to make high-confidence predictions even on OOD inputs, demonstrating their inability to recognize data outside the training distribution as OOD [6, 7]. The reliability of OOD detection is especially critical in applications like healthcare and autonomous driving, where overconfident predictions on unfamiliar data could have serious consequences [8, 9].
@@
Figure 1: The challenge of spurious correlations in OOD detection. ID classes (dog, fox, wolf) appear in correlated backgrounds (grass, autumn, snow), with majority groups relying on context shortcuts (blue frames). SP-OOD samples share these contexts, complicating detection. Non-spurious OOD samples differ in context and lack both spurious and core features.
Recent efforts to mitigate spurious correlations in OOD detection can be grouped into several broad categories. Outlier Exposure (OE) techniques reduce reliance on spurious correlations by incorporating synthetic OOD samples in their training [21, 22, 23, 24]. Other methods focus on modifying training objectives to explicitly discourage models from depending on spurious features [25, 26]. However, these methods typically require extensive training and the generation of OOD data. In contrast, several post-hoc approaches have been proposed that bypass the limitations of training-heavy approaches and offer fast and light alternatives [27, 28, 29, 30, 31, 32]. Despite their promise, these methods are often evaluated on limited synthetic datasets and backbones, or rely on multiple modalities.
@@
  • We propose SPROD, a post-hoc OOD detection method that directly addresses unknown spurious correlations by design and outperforms the state-of-the-art.
@@ 2 Related Work
OOD detection methods can be categorized into training-time and post-hoc approaches [37]. Training-time methods leverage auxiliary OOD samples (Outlier Exposure) [12, 38, 39] or apply regularization [40, 41, 42, 43] to enhance OOD detection. Post-hoc methods, in contrast, derive OOD scores from base classifiers without modifying training [37]. Overall, post-hoc methods offer simplicity and competitive performance [37], making them practical under limited data or training resources.
Among post-hoc methods, several approaches apply transformations to model logits to derive OOD scores. MSP [7] uses the maximum softmax probability. The energy-based method [13] uses the log-sum-exp of logits. MLS [44] uses the maximum logit and introduces KL Matching (KLM), which computes the KL divergence between the softmax output and class-specific reference distributions.
Another class of post-hoc methods uses feature-space distances for OOD detection. MDS [11] fits class-conditional Gaussians to pre-logit features and computes Mahalanobis distance. RMDS [45] refines MDS by modeling background distributions with an unconditional Gaussian on ID data. KNN [46] uses distances to nearest ID training samples. SHE [47] stores ID features as templates and scores by their distance to test samples.
A few methods combine information from both feature and logit spaces. ReAct [48] thresholds features by activation magnitude before applying scoring functions like energy. ViM [49] adds a virtual logit from the residual norm between input features and the ID subspace, then applies softmax over the extended logits. Some methods also exploit gradient space for OOD scoring [50, 51]. GradNorm [50] computes the KL divergence to a uniform distribution and uses the gradient norm (w.r.t. the penultimate layer) as the score.
@@
Recent advances in vision-language models [53] have led to a category of zero-shot OOD detection methods [29, 30, 31, 32] that use textual inputs, such as class names or attribute descriptions, to define ID data and identify OOD samples. While some report results on Waterbirds SP-OOD, this is not their main focus. Moreover, lacking training data with spurious correlations, their zero-shot setting does not fully capture the spurious OOD challenge. In contrast to these approaches, which rely on explicit text for OOD scoring, our method refines visual prototypes using only ID features and class labels from the training set.
@@ 3.1 Problem Setup
This paper addresses Out-of-Distribution (OOD) detection under spurious correlations in a supervised classification setting. Let $\mathcal{X}$ be the input space and $\mathcal{Y}={1,\dots,C}$ the label set. The in-distribution (ID) training data $\mathcal{D}_{\text{in}}=\{(x_{i},y_{i})\}_{i=1}^{N}$ comprises samples from a joint distribution $P_{\mathcal{X},\mathcal{Y}}$. A neural network $f_{\theta}$ maps each input $x_{i}$ to a feature embedding $h_{i}=f_{\theta}(x_{i})$. This network is typically pretrained or fine-tuned on the training data. OOD detection identifies test samples from distributions outside the training data, including unseen classes with labels beyond the training space.
@@ 3.2 Score Calculation
A key element in OOD detection is the scoring function $\mathcal{S}(x)$, which maps an input to a scalar indicating its similarity to the in-distribution (ID) training data. This score is typically derived from a model’s learned representations or its predictive outputs. OOD detection is performed based on this score function. An effective OOD method offers distinct and well-separated distributions of scores for ID and OOD samples.
Prior studies show that feature-based scores, derived from intermediate representations [10, 13, 46], typically outperform those based on output probabilities [7], especially in the presence of spurious correlations [14]. Our experiments further support this trend in certain settings, suggesting that output-based methods (especially those relying on softmax probabilities) may be less reliable when model confidence is influenced by spurious features. To better understand this difference, consider two probabilistic perspectives in classification: discriminative models directly estimate $p(y|x)$ and focus on separating classes, while generative models estimate $p(x|y)$ and capture how data $x$ is distributed for each class $y$. Most of distance-based OOD detection methods can be viewed as approximating a generative approach in feature space.
While directly modeling $p(x|y)$ is generally effective for classifying in-distribution samples, its utility for OOD detection can be limited, particularly when models are trained solely on ID data without exposure to OOD examples. The softmax function forces a normalized probability distribution over the known classes, which can degrade OOD detection performance. For instance, an OOD sample that lies far from all class distributions may receive a high softmax probability if it is only slightly distant from a decision boundary. Conversely, an in-distribution sample situated near the boundary between multiple classes could receive a low maximum softmax probability, potentially leading to its misclassification as OOD. In contrast, distance-based approaches, which rely on the class-conditional distribution $p(x|y)$, are by nature more robust in these scenarios. OOD samples that share few characteristics with any known class typically exhibit low likelihood under all class-conditional distributions $p(x|y)$, and can be reliably identified, regardless of their proximity to decision boundaries.
@@ 4.1 Stage 1: Initial Prototype Construction
Given a pretrained feature extractor $f_{\theta}$, we first obtain feature embeddings $h_{i}=f_{\theta}(x_{i})$ for each training sample $x_{i}$. To ensure uniformity in feature representation, these embeddings are normalized to have unit norm $z_{i}=h_{i}/\lVert h_{i}\rVert_{2}$. For each in-distribution class $c\in\mathcal{Y}$, an initial prototype $\mathbf{p}_{c}$ is computed as the mean of these normalized embeddings: $\mathbf{p}_{c}=1/N_{c}\sum_{i:y_{i}=c}z_{i}$, where $N_{c}$ is the number of samples in class $c$. Each $\mathbf{p}_{c}$ serves as an initial estimate of the class centroid in the normalized feature space. A query sample $x_{q}$ (with normalized embedding $z_{q}$) is typically classified to the class $c$ whose prototype $\mathbf{p}_{c}$ is the closest (e.g., using Euclidean distance $d(z_{q},\mathbf{p}_{c})$). We use $d(z_{q},\mathbf{p}_{c})$ as the scoring function for OOD detection. While our empirical results demonstrate the effectiveness of the naïve prototypical method in both SP-OOD and NSP-OOD settings, this approach remains vulnerable to biases from spurious correlations in the training data. As a result, the prototypes become skewed toward majority groups within each class (see part I of Figure 3). This bias leads to a scenario where SP-OOD samples (represented by purple squares in Figure 3) may be erroneously classified as in-distribution due to their proximity to these biased prototypes. These limitations motivate the subsequent refinement stages.
@@ 4.3 Stage 3: Group Prototype Refinement
The refined prototypes $\mathbf{p}_{c}^{\text{maj}}$ and $\mathbf{p}_{c\rightarrow m}^{\text{min}}$, further enhance bias reduction by the proposed refitting process over unknown groups of each class. During classification and OOD detection, the query embedding $z_{q}$ is compared to all group-specific prototypes, and the final prediction is based on the nearest prototype, regardless of group type. This multiple-prototype approach reduces the likelihood of OOD samples being erroneously classified due to shared spurious attributes with any single prototype (see part III of Figure 3). For each sample, OOD score is simply calculated based on the distance to the nearest group prototype. The overall procedure of our prototypical-based OOD detection method is outlined in the Algorithm 1.
@@
11:       Separate samples into majority $\mathcal{S}_{c}^{\text{maj}}$ and minority $\mathcal{S}_{c\rightarrow m}^{\text{min}}$ based on proximity to $\mathbf{p}_{c}^{\text{corr}}$ or $\mathbf{p}_{c\rightarrow m}^{\text{misc}}$
@@ 5.1 Experimental Setup
We evaluate our method against a comprehensive suite of 11 post-hoc OOD detection approaches: MSP [7], MDS [11], RMDS [45], Energy [13], GradNorm [50], ReAct [48], MaxLogit [44], KLM [44], VIM [49], KNN [46], and SHE [47]. We assess performance primarily using the Area Under the Receiver Operating Characteristic curve (AUROC), a threshold-independent metric, and the False Positive Rate at 95% True Positive Rate (FPR@95). Additional metrics, including the Area Under the Precision-Recall curve (AUPR), are given in the supplementary material. We repeat all experiments five times with different random seeds and report the mean and standard deviation.
In this experiments section we focus on the more challenging SP-OOD scenarios, particularly those with the highest degree of spurious correlation in each dataset, as our method achieves near-perfect performance on NSP-OOD samples. Results for less challenging settings, along with a broader analysis involving a variety of transformer-based and convolutional backbones are included in the supplementary material. For consistency, the results in this section primarily use the widely-adopted ResNet-50 [55] backbone (with ResNet-18 additionally used in one analysis).
@@ 5.2 Datasets
For evaluating SP-OOD detection, we utilize 5 diverse datasets, whose properties are summarized in Table 1. Additional details and examples for all datasets, including the NSP datasets and their evaluation results, are included in the supplementary material.
The datasets include established datasets for OOD detection, namely Waterbirds (WB) [33] and CelebA (CA) [34]. While widely adopted, these datasets present limitations in terms of scale and realism (CelebA, in particular, is noted for its label noise [57]), making them insufficient for comprehensive evaluation. To address this, our evaluation incorporates three additional datasets designed to test robustness under diverse conditions, including multi-class scenarios, multiple spurious attributes, and real-world complexity. First, UrbanCars (UC) [35] is a binary classification dataset (urban vs. country cars) with two spurious attributes: background and a co-occurring object, both correlated with the class, making it a challenging multi-spurious benchmark. Next, we introduce Animals MetaCoCo (AMC), a realistic SP-OOD benchmark created by subsampling and curating animal categories from MetaCoCo [58]. It contains 26 distinct classes, each with 8 different background types serving as imbalance shortcut attributes. We use a class-level leave-2-out setting, where two classes are held out as SP-OOD and the rest are treated as ID. The significant similarity in background distributions between the ID and OOD splits makes Animals MetaCoCo a particularly challenging multi-class benchmark for SP-OOD detection (Figure 1). Last, we also use the Spurious ImageNet dataset introduced in [36] as SP-OOD data. This dataset consists of real-world images that contain only spurious features, such as bird feeders or graffiti, without the actual class objects, yet are consistently misclassified as specific ImageNet classes. These OOD samples are constructed for a subset of 100 ImageNet classes identified to rely on harmful spurious correlations. We refer to this subset of classes as Sp-ImageNet100 (SpI), a name we use for clarity, and treat it as the ID data.

### v2

@@ Abstract
Out-of-distribution (OOD) detection is crucial for ensuring the reliability and safety of machine learning models in real-world applications, where they frequently face data distributions unseen during training. Despite progress, existing methods are often vulnerable to spurious correlations that mislead models and compromise robustness. To address this, we propose SPROD, a novel prototype-based OOD detection approach that explicitly addresses the challenge posed by unknown spurious correlations. Our post-hoc method refines class prototypes to mitigate bias from spurious features without additional data or hyperparameter tuning, and is broadly applicable across diverse backbones and OOD detection settings. We conduct a comprehensive spurious correlation OOD detection benchmarking, comparing our method against existing approaches and demonstrating its superior performance across challenging OOD datasets, such as CelebA, Waterbirds, UrbanCars, Spurious Imagenet, and the newly introduced Animals MetaCoCo. On average, SPROD improves AUROC by 4.8% and FPR@95 by 9.4% over the second best.
@@ 1 Introduction
Machine learning systems in real-world applications often encounter out-of-distribution (OOD) inputs, which are samples from distributions different from the training data. These inputs require cautious handling to prevent overconfident mispredictions during inference [1]. This makes OOD detection crucial, as it aims to identify whether an input belongs to the known distribution or not. Yet, deep neural networks, widely used in vision tasks [2, 3, 4, 5], tend to make high-confidence predictions even on OOD inputs, demonstrating their inability to recognize data outside the training distribution as OOD [6, 7]. The reliability of OOD detection is especially critical in applications like healthcare and autonomous driving, where overconfident predictions on unfamiliar data could have serious consequences [8, 9].
@@
Figure 1: The challenge of spurious correlations in OOD detection. ID classes (dog, fox, wolf) appear in correlated backgrounds (grass, autumn, snow), with majority groups relying on context shortcuts (blue frames). SP-OOD samples share the same contextual backgrounds, making detection more difficult. NSP-OOD samples differ in context and lack both spurious and core features.
Recent efforts to mitigate spurious correlations in OOD detection can be grouped into several broad categories. Outlier Exposure (OE) techniques reduce reliance on spurious correlations by incorporating synthetic OOD samples in their training [21, 22, 23, 24]. Other methods focus on modifying training objectives to explicitly discourage models from depending on spurious features [25, 26]. However, these methods usually require retraining and generating additional OOD data. In contrast, several post-hoc approaches have been proposed that bypass the limitations of training-heavy approaches and offer fast and light alternatives [27, 28, 29, 30, 31, 32]. Despite their promise, these methods are often tested on limited synthetic datasets or specific backbones, and some rely on multiple modalities.
@@
  • We propose SPROD, a post-hoc OOD detection method that directly addresses unknown spurious correlations by design and outperforms the state-of-the-art. We provide theoretical insight into how the proposed method mitigates spurious bias.
@@ 2 Related Work
OOD detection methods can be categorized into training-time and post-hoc approaches [37]. Training-time methods leverage auxiliary OOD samples (Outlier Exposure) [12, 38, 39] or apply regularization [40, 41, 42, 43] to enhance OOD detection. Post-hoc methods, in contrast, derive OOD scores from base classifiers without modifying training [37]. Overall, post-hoc methods offer simplicity and competitive performance [37], making them practical under limited data or training resources. Among post-hoc methods, several approaches apply transformations to model logits to derive OOD scores. MSP [7] uses the maximum softmax probability, the energy-based method [13] computes the log-sum-exp of logits, MLS [44] uses the maximum logit and introduces KL Matching (KLM) based on KL divergence, and GEN [45] employs generalized entropy of softmax outputs.
Another class of post-hoc methods detects OOD samples via feature-space distances. MDS [11] fits class-conditional Gaussians to pre-logit features and computes Mahalanobis distances, refined by RMDS [46] with an unconditional Gaussian on ID data. KNN [47] uses distances to nearest ID samples. SHE [48] scores samples by their distance to stored ID feature templates. NNGuide [49] leverages nearest-neighbor guidance to adjust test features toward the ID manifold. Relation [50] constructs a graph over training embeddings and detects outliers via relational anomalies. NECO [51] scores samples by their feature alignment with class weight vectors, leveraging neural collapse geometry. SCALE [52] separates ID and OOD samples by scaling penultimate-layer activations. FDBD [53] measures features’ regularized mean distance to the classifier’s decision boundaries. NCI [54] scores samples by their distance to class weight vectors, filtered by feature norms.
Prototype-based methods shape class representations for distance-based OOD scoring. Classical approaches like MDS [11] and its variants [46] are closely related, as they model each class by a centroid in feature space, optionally using class-conditional covariances to compute Mahalanobis distances. Recent works extend this via explicit training objectives. CIDER [55] learns hyperspherical embeddings by jointly enforcing intra-class compactness and inter-class dispersion, thereby improving ID and OOD separability. PALM [56] represents each class as a mixture of learnable prototypes and optimizes a maximum-likelihood and contrastive objective, updating prototypes and backbone features jointly during training. PROWL [57] also leverages prototype representations, but for pixel-level OOD detection in segmentation. While these methods share a prototypical framework, SPROD differs as a post-hoc method operating on pretrained backbones and is explicitly designed to mitigate the negative effects of unknown spurious correlations. Appendix H further analyzes a variant, SPROD-KMeans, which connects to mixture-of-prototypes ideas while remaining fully post-hoc.
A few methods combine information from both feature and logit spaces. ReAct [58] thresholds activations before applying energy-based scoring. ViM [59] adds a virtual logit from the residual norm between input features and the ID subspace and applies softmax over extended logits. ASH [60] prunes high-magnitude activations and rescales remaining features before logit computation, improving energy-based OOD separability. Some methods also exploit gradient space for OOD scoring [61, 62]. GradNorm [61] computes the KL divergence to a uniform distribution and uses the gradient norm (w.r.t. the penultimate layer) as the score.
@@
Recent advances in vision-language models [64] have led to a category of zero-shot OOD detection methods [29, 30, 31, 32] that use textual inputs, such as class names or attribute descriptions, to define ID data and identify OOD samples. While some report results on Waterbirds SP-OOD, this is not their main focus. Moreover, lacking training data with spurious correlations, their zero-shot setting does not fully capture the SP-OOD challenge. In contrast to these approaches, which rely on explicit text for OOD scoring, our method refines visual prototypes using only ID features and class labels from the training set. A detailed review of studies regarding SP-OOD can be found in Appendix B.
@@ 3.1 Problem Setup
This paper addresses Out-of-Distribution (OOD) detection under spurious correlations in a supervised classification setting. Let $\mathcal{X}$ be the input space and $\mathcal{Y}=\{1,\dots,C\}$ the label set. The in-distribution (ID) training data $\mathcal{D}_{\text{in}}=\{(x_{i},y_{i})\}_{i=1}^{N}$ comprises samples from a joint distribution $P_{\mathcal{X},\mathcal{Y}}$. A neural network $f_{\theta}$ maps each input $x_{i}$ to a feature embedding $h_{i}=f_{\theta}(x_{i})$. This network is typically pretrained or fine-tuned on the training data. OOD detection aims to identify test samples from distributions not seen during training, including those from unseen classes.
@@ 3.2 Score Calculation
A key element in OOD detection is the scoring function $\mathcal{S}(x)$, which assigns a scalar value reflecting how likely an input belongs to the in-distribution (ID). This score is typically derived from a model’s learned representations or its predictive outputs. OOD detection is performed based on this score function. An effective OOD method offers distinct and well-separated distributions of scores for ID and OOD samples.
Prior studies show that feature-based scores, derived from intermediate representations [10, 13, 47], typically outperform those based on output probabilities [7], especially in the presence of spurious correlations [14]. Our experiments further support this trend in certain settings, suggesting that output-based methods (especially those relying on softmax probabilities) may be less reliable when model confidence is influenced by spurious features. To better understand this difference, consider two probabilistic perspectives in classification: discriminative models directly estimate $p(y|x)$ and focus on separating classes, while generative models estimate $p(x|y)$ and capture how data $x$ is distributed for each class $y$. Most distance-based OOD detection methods can be viewed as approximating a generative approach in feature space.
While directly modeling $p(y|x)$ is generally effective for classifying ID samples in standard settings, discriminative approaches have been reported to be more sensitive under distribution shifts, such as in continual learning [65] or in the presence of spurious correlations [66]. Furthermore, their utility for OOD detection can be limited, particularly when methods are optimized solely on ID data without exposure to OOD examples. The softmax function forces a normalized probability distribution over the known classes, which can degrade OOD detection performance. For instance, an OOD sample that lies far from all class distributions may receive a high softmax probability if it is only slightly distant from a decision boundary. Conversely, an ID sample situated near the boundary between multiple classes could receive a low maximum softmax probability, potentially leading to its misclassification as OOD. In contrast, distance-based approaches, which rely on the class-conditional distribution $p(x|y)$, are by nature more robust in these scenarios. OOD samples that share few characteristics with any known class typically exhibit low likelihood under all class-conditional distributions $p(x|y)$, and can be reliably identified, regardless of their proximity to decision boundaries.
@@ 4.1 Stage 1: Initial Prototype Construction
Given a pretrained feature extractor $f_{\theta}$, we first obtain feature embeddings $h_{i}=f_{\theta}(x_{i})$ for each training sample $x_{i}$. To ensure uniformity in feature representation, these embeddings are normalized to have unit norm $z_{i}=h_{i}/\lVert h_{i}\rVert_{2}$. For each ID class $c\in\mathcal{Y}$, an initial prototype $\mathbf{p}_{c}$ is computed as the mean of these normalized embeddings: $\mathbf{p}_{c}=1/N_{c}\sum_{i:y_{i}=c}z_{i}$, where $N_{c}$ is the number of samples in class $c$. Each $\mathbf{p}_{c}$ serves as an initial estimate of the class centroid in the normalized feature space. A query sample $x_{q}$ (with normalized embedding $z_{q}$) is typically classified to the class $c$ whose prototype $\mathbf{p}_{c}$ is the closest (e.g., using Euclidean distance $d(z_{q},\mathbf{p}_{c})$). We use $d(z_{q},\mathbf{p}_{c})$ as the scoring function for OOD detection. While our empirical results demonstrate the effectiveness of the naïve prototypical method in both SP-OOD and NSP-OOD settings, this approach remains vulnerable to biases from spurious correlations in the training data. As a result, the prototypes become skewed toward majority groups within each class (see part I of Figure 3). This bias leads to a scenario where SP-OOD samples (represented by purple squares in Figure 3) may be erroneously classified as ID due to their proximity to these biased prototypes. These limitations motivate the subsequent refinement stages.
@@ 4.3 Stage 3: Group Prototype Refinement
The refined prototypes $\mathbf{p}_{c}^{\text{maj}}$ and $\mathbf{p}_{c\rightarrow m}^{\text{min}}$, further reduce bias through the proposed refitting process. During classification and OOD detection, the query embedding $z_{q}$ is compared to all group-specific prototypes, and the final prediction is based on the nearest prototype, regardless of group type. This multiple-prototype approach reduces the likelihood of OOD samples being erroneously classified due to shared spurious attributes with any single prototype (see part III of Figure 3). For each sample, the OOD score is simply calculated based on the distance to the nearest group prototype.
@@
11:    Construct majority $\mathcal{S}_{c}^{\text{maj}}$ and minority $\{\mathcal{S}_{c\rightarrow m}^{\text{min}}\}_{m\neq c}$ groups based on proximity to $\mathbf{p}_{c}^{\text{corr}}$ or $\{\mathbf{p}_{c\rightarrow m}^{\text{misc}}\}_{m\neq c}$
@@
The overall procedure of our prototype refinement strategy is outlined in Algorithm 1 and a theoretical justification of how the proposed procedure mitigates spurious bias is provided in Appendix A.
@@ 5.1 Experimental Setup
We evaluate our method against a comprehensive suite of 19 post-hoc OOD detection approaches: MSP [7], MDS [11], RMDS [46], Energy [13], GradNorm [61], ReAct [58], MaxLogit [44], MLS & KLM [44], VIM [59], KNN [47], SHE [48], ASH [60], NECO [51], NNGuide [49], Relation [50], SCALE [52], fDBD [53], and NCI [54]. We assess performance primarily using the Area Under the Receiver Operating Characteristic curve (AUROC), a threshold-independent metric, and the False Positive Rate at 95% True Positive Rate (FPR@95). Additional metrics, including the Area Under the Precision-Recall curve (AUPR), are provided in Appendix G. We repeat all experiments five times with different random seeds and report the mean and standard deviation.
In this experiments section, we focus on the more challenging SP-OOD scenarios, particularly those with the highest degree of spurious correlation in each dataset, as our method achieves near-perfect performance on far NSP-OOD samples. Detailed results for the NSP-OOD setting are provided in Appendix E. Further analyses of SPROD’s stages, along with broader evaluations across various transformer-based and convolutional backbones, are included in Appendix F and Appendix G. For consistency, the results in this section primarily use the widely adopted ResNet-50 [68] backbone, with ResNet-18 additionally used in one analysis. Beyond SP-OOD benchmarks, we also evaluate our approach under conventional (standard) OOD settings to further demonstrate its general applicability.
@@ 5.2 Datasets
For evaluating SP-OOD detection, we utilize five diverse datasets, whose properties are summarized in Table 1. Additional details and visual examples for all datasets, including the NSP datasets and their evaluation setup, are provided in Appendix D. The datasets include Waterbirds (WB) [33] and CelebA (CA) [34]. While widely adopted, these datasets present limitations in terms of scale and realism (CelebA, in particular, is noted for its label noise [70]), making them insufficient for comprehensive evaluation. To address this, our evaluation incorporates three additional datasets designed to test robustness under diverse conditions, including multi-class scenarios, multiple spurious attributes, and real-world complexity.
UrbanCars (UC) [35] is a binary classification dataset (urban vs. country cars) with two spurious attributes: background and a co-occurring object, both correlated with the class, making it a challenging multi-spurious benchmark. Next, we introduce Animals MetaCoCo (AMC), a realistic SP-OOD benchmark created by subsampling and curating animal categories from MetaCoCo [71]. It contains 26 distinct classes, each with 8 different background types serving as imbalance shortcut attributes. We use a class-level leave-2-out setting, where two classes are held out as SP-OOD and the rest are treated as ID. The significant similarity in background distributions between the ID and OOD splits makes Animals MetaCoCo a particularly challenging multi-class benchmark for SP-OOD detection (Figure 1). Last, we also use the Spurious ImageNet dataset introduced in [36] as SP-OOD data. This dataset consists of real-world images that contain only spurious features, such as bird feeders or graffiti, without the actual class objects, yet are consistently misclassified as specific ImageNet classes. These OOD samples are constructed for a subset of 100 ImageNet classes identified to rely on harmful spurious correlations. We refer to this subset of classes as Sp-ImageNet100 (SpI), a name we use for clarity, and treat it as the ID data.
For conventional OOD evaluations, we follow the same settings as in OpenOOD [37]. Specifically, we use CIFAR-10 [72], CIFAR-100 [72], and ImageNet-1k [73], along with their respective near-OOD and far-OOD datasets, to ensure consistency with widely adopted benchmarks in the field.